# CutPaste&Find: Efficient Multimodal Hallucination Detector with Visual-aid Knowledge Base

### v1

@@ 1 Introduction
"However, these methods have several limitations. First, they rely heavily on external API calls. As reported in Table 1, Woodpecker and LogicCheckGPT make 4-6 API calls to detect each sentence. This dependence also limits offline usability, making deployment infeasible in scenarios with restricted API access or privacy concerns. Second, they require iterative interactions with LVLMs to detect hallucinations. For instance, LogicCheckGPT generates multiple queries and performs consistency checks across responses. This iterative process not only increases latency but also amplifies token consumption, making it impractical for large-scale applications or real-time inference.
@@ 2.1 LVLMs
With the rapid advancement of large language models (LLMs) Ouyang et al. (2022); Zhao et al. (2023); Pan et al. (2024); Wu et al. (2024e, f, d, c), integrating their general intelligence into multimodal domains has garnered significant interest. This has led to the emergence of large vision-language models (LVLMs) Ye et al. (2023); Zhu et al. (2023); Liu et al. (2023a); Li et al. (2023a); Dai et al. (2023a); Bai et al. (2023a), designed to understand and generate multimodal content under instructions. Most LVLMs adopt a two-stage training paradigm: multimodal alignment pre-training followed by instruction tuning, where an alignment module processes multimodal inputs before passing them to an LLM for response generation. For instance, mPLUG-Owl Ye et al. (2023) pre-trains both the encoder and alignment module before fine-tuning LLaMa Touvron et al. (2023) with low-rank adaptation. LLaVA Liu et al. (2023a) pre-trains the alignment network and fine-tunes it alongside Vicuna Chiang et al. (2023) on constructed instructions. In contrast, MiniGPT-4 Zhu et al. (2023) only fine-tunes the cross-modal alignment network while keeping other components frozen.
@@ 3.3.2 Visual verification with Visual Similarity
To complement VQA, we validate hallucination using Visual-aid Knowledge Base. For an entity-attribute pair ($a_{i}$,$e_{i}$), we retrieve a list of representative images $\hat{C}_{i}$ from the datastore and compute visual representations $f_{i}$ (for the cropped image) and $\hat{F}_{i}$ (for retrieved images) using a pre-trained visual encoder (e.g., CLIP). The similarity score $s^{v}_{i}$ is calculated between $f_{i}$ and $\hat{F}_{i}$; lower scores indicate higher likelihoods of hallucination. A similar approach is applied to validate relationships $s_{i},r_{i},o_{i}$.

### v2

@@ 1 Introduction
"However, these methods have several limitations. First, they rely heavily on external API calls. As reported in Table˜1, Woodpecker and LogicCheckGPT make 4-6 API calls to detect each sentence. This dependence also limits offline usability, making deployment infeasible in scenarios with restricted API access or privacy concerns. Second, they require iterative interactions with LVLMs to detect hallucinations. For instance, LogicCheckGPT generates multiple queries and performs consistency checks across responses. This iterative process not only increases latency but also amplifies token consumption, making it impractical for large-scale applications or real-time inference.
@@ 2.1 LVLMs
With the rapid advancement of large language models (LLMs) Ouyang et al. (2022); Zhao et al. (2023); Pan et al. (2024); Wu et al. (2024g, h, e, d), integrating their general intelligence into multimodal domains has garnered significant interest. This has led to the emergence of large vision-language models (LVLMs) Ye et al. (2023); Zhu et al. (2023); Liu et al. (2023a); Li et al. (2023a); Dai et al. (2023a); Bai et al. (2023a), designed to understand and generate multimodal content under instructions. Most LVLMs adopt a two-stage training paradigm: multimodal alignment pre-training followed by instruction tuning, where an alignment module processes multimodal inputs before passing them to an LLM for response generation. For instance, mPLUG-Owl Ye et al. (2023) pre-trains both the encoder and alignment module before fine-tuning LLaMa Touvron et al. (2023) with low-rank adaptation. LLaVA Liu et al. (2023a) pre-trains the alignment network and fine-tunes it alongside Vicuna Chiang et al. (2023) on constructed instructions. In contrast, MiniGPT-4 Zhu et al. (2023) only fine-tunes the cross-modal alignment network while keeping other components frozen.
@@ 2.2 Hallucination in LVLMs
Complementary to these methods, a growing body of research aims to reduce hallucination at its semantic representation level. Works from Nguyen et al. (2021) and Nguyen et al. (2024c) demonstrate that explicit global semantics control helps anchor generations to context. In a similar vein, Nguyen and Luu (2021); Nguyen et al. (2024d); Nguyen and Nguyen (2024) show that refining latent topic structures can improve factual consistency in text generation.
In multimodal settings, angular margin-based contrastive learning (Nguyen et al., 2024a, 2022) has been effective in tightening inter-modal alignment, which is crucial for grounding. Likewise, models like (Vo et al., 2024) and (Nguyen et al., 2023a) enhance grounding via stronger joint embeddings or partial alignment. These representation-level improvements are vital for hallucination prevention, especially when training or decoding resources are limited.
Hallucination is further exacerbated in video-language models, where temporal reasoning adds complexity. Models like DemaFormer (Nguyen et al., 2023b), READ (Nguyen et al., 2024b) introduce novel architectures or adapters to better model temporal grounding. Motion-aware and multi-scale contrastive learning further improve alignment between textual queries and dynamic visual content—reducing hallucinations by anchoring generations in fine-grained temporal cues.
Finally, advances in neural topic modeling (e.g., (Wu et al., 2024e, 2023a, d, f, c, 2023b) contribute insights into how interpretable, structured representations can guide generation models. This direction suggests that structured semantic scaffolding—via topic modeling or contrastive optimization—may offer long-term solutions to hallucination by grounding language in disentangled conceptual spaces.
@@ 3.3.2 Visual verification with Visual Similarity
To complement VQA, we validate hallucination using Visual-aid Knowledge Base. For an entity-attribute pair ($a_{i}$$e_{i}$), we retrieve a list of representative images $\hat{C}_{i}$ from the datastore and compute visual representations $f_{i}$ (for the cropped image) and $\hat{F}_{i}$ (for retrieved images) using a pre-trained visual encoder (e.g., CLIP). The similarity score $s^{v}_{i}$ is calculated between $f_{i}$ and $\hat{F}_{i}$; lower scores indicate higher likelihoods of hallucination. A similar approach is applied to validate relationships $s_{i},r_{i},o_{i}$.